OpenStudy (anonymous):

A box contains 15 cards. There are 5 red cards labeled 1, 2, 3, 4, and 5. There are 5 blue cards labeled 1, 2, 3, 4, and 5. There are 5 green cards labeled 1, 2, 3, 4, and 5. One card is drawn from the box. What is the probability that it is a green card or it is a red card with a 5? A. 2/3 B. 1/45 C. 2/5 D. 1/3 Medal

OpenStudy (anonymous):

@Phebe

OpenStudy (sparrow2):

it's c

OpenStudy (anonymous):

\[\frac{5}{15}+\frac{1}{15}=\frac{2}{5} \]

OpenStudy (sparrow2):